OpenStudy (anonymous):

can anyone help me solve 2 problems with substitution? :)

OpenStudy (anonymous):

2 = 2y - x 23 = 5y - 4x

OpenStudy (rogue):

If you rearrange the first equation, you can get: x = 2y - 2 You can then plug that into the x in the second equation. 23 = 5y - 4(2y - 2) Then just solve that for y. After you get y, plug y back into x = 2y - 2 to get your x value. After you have both the x and y values, your done!

OpenStudy (anonymous):

thank you! i got -12, -5

OpenStudy (rogue):

Good job :)

OpenStudy (anonymous):

& i also have a last problem like that

OpenStudy (anonymous):

if you have time:)

OpenStudy (rogue):

Sure, hit me.

OpenStudy (anonymous):

The measure of one acute angle in a right triangle is four times the measure of the acute angle. Write and solve a system of equations to find the measure of the acute angles.

OpenStudy (rogue):

Alright, so what do you know about the total degree measure of a triangle? The degree measure of a right angle?

OpenStudy (anonymous):

180 & 90

OpenStudy (rogue):

Alright, good. Total angle measure = angle 1 + angle 2 + angle 3. 180 = 90 + angle 2 + angle 3 90 = angle 2 + angle 3 So the sum of the two acute angles is 90 degrees.

OpenStudy (anonymous):

yess:)

OpenStudy (rogue):

If you let x = measure of the smaller acute angle, what is the measure of the bigger acute angle?

OpenStudy (anonymous):

that's the part i don't understand.

OpenStudy (rogue):

Well, the bigger acute angle is 4 times as much as the smaller one. Small acute angle = x Big acute angle = 4x

OpenStudy (rogue):

And their sum is 90 degrees, so 90 = x + 4x

OpenStudy (anonymous):

ohh!

OpenStudy (rogue):

You can do the rest right? :)

OpenStudy (anonymous):

yes thank you!! :) i wish i had your brain lol

OpenStudy (rogue):

Haha, thanks, don't worry, you'll get there :)

OpenStudy (anonymous):

and one more question, but short

OpenStudy (anonymous):

x + 3 = y 3x + 4y = 7 so that would be; 3x + 4(x + 3) = 7 ?

OpenStudy (rogue):

Yep, nice :)

OpenStudy (anonymous):

thanks! :)

OpenStudy (rogue):

For systems, just try to find a way to substitute one equation into the other.

OpenStudy (anonymous):

Okay! thank you:D

OpenStudy (rogue):

Good luck with the rest, byes. :)

OpenStudy (anonymous):

byee:)